OpenStudy (anonymous):

Help? Please and thank you! What is the y coordinate of the y−intercept of the line that passes through the points (−3, −1) and (1, 7)?

OpenStudy (anonymous):

-3 -1 1 7

OpenStudy (anonymous):

You have to find the slope of the line!

OpenStudy (anonymous):

Slope is...\[\frac{ RISE }{ RUN }\]

OpenStudy (anonymous):

Lagging is my life.

OpenStudy (anonymous):

would that be equal to 1/2 for slope? or no?

OpenStudy (anonymous):

Right:)

OpenStudy (anonymous):

Well, what is my question asking for? the y coordinate?

OpenStudy (anonymous):

Little confused.-.

OpenStudy (anonymous):

Yes! To find where the line crosses the Y axes!

OpenStudy (anonymous):

-2, 0? Random guess. Lol

OpenStudy (anonymous):

\[y – y _{1} = m(x – x _{1})\]

OpenStudy (anonymous):

Now using that formula I gave you plug in one of the Points that are given and solve for y.

OpenStudy (anonymous):

-1 + 7 = m(-3 + 1)

OpenStudy (anonymous):

- 7*

OpenStudy (anonymous):

Not exactly...Only use one set of points and plug it into the y and x with a tiny 1...

OpenStudy (anonymous):

More like... y+1= M(x+3)

OpenStudy (anonymous):

And M means slope so plug that in too!

OpenStudy (anonymous):

y + 1 = 1/2(x + 3)

OpenStudy (anonymous):

Yup, now solve for Y.

OpenStudy (anonymous):

y = 1/2x + 1/2 ?

OpenStudy (anonymous):

Yes:)

OpenStudy (anonymous):

This form is call slope intercept form. To find y plug in 0 for X.

OpenStudy (anonymous):

y = 1/2(0) + 1/2 y = 1/2 ?

OpenStudy (anonymous):

(0,1/2)

OpenStudy (anonymous):

Now, this will be tricky to put in the written response, because fractions turn from 1/2 to 12. Lol. Should I just put 0.5?

OpenStudy (anonymous):

Wait...before you write it lets check it...cause something doesnt seem right...

OpenStudy (anonymous):

I feel like it should be a whole number, my teacher said it should be for these ones.

OpenStudy (anonymous):

I think is should be (0,5)...

OpenStudy (anonymous):

@mathstudent55 Maybe they know if we did anything wrong.

OpenStudy (anonymous):

I figured out!

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

The slope is 2!

OpenStudy (anonymous):

we have to redo that with the slope of 2.

OpenStudy (anonymous):

Because its Y/X not X/Y.

OpenStudy (anonymous):

OpenStudy is lagging for me.

OpenStudy (anonymous):

Its okay! it is (0,5)

OpenStudy (anonymous):

\[y+1=2(x+3)\]

OpenStudy (anonymous):

\[y=2x+5\]

OpenStudy (anonymous):

@Luigi0210 I really need to know if he's right. I'm so confused.

OpenStudy (anonymous):

Sorry for confusing you...:(

OpenStudy (anonymous):

http://www.webmath.com/_answer.php Its right, check...

OpenStudy (anonymous):

Only if OpenStudy was a real person...

OpenStudy (luigi0210):

So you're just looking for the y-coordinate of the y intercept?

OpenStudy (anonymous):

Yes.

OpenStudy (luigi0210):

Well you know that the y-intercept is when x=0 right?

OpenStudy (anonymous):

Yes, but now I don't even know what the slope is anymore. 1/2 or 2 or 5 lol.

OpenStudy (luigi0210):

Slope is: \[m=\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

-3, -1 1, 7

OpenStudy (anonymous):

-1 - 7 ------ -3 - 1 ?

OpenStudy (luigi0210):

Just plug them in and solve. and either way works, you get the same answer

OpenStudy (anonymous):

-8 -- -4 ?

OpenStudy (luigi0210):

Yup, making the slope 2. Now we plug that information into the point-slope formula: y-y1=m(x-x1)

OpenStudy (luigi0210):

Open was right so far, except for the slope

OpenStudy (anonymous):

y - 1 = 2(x - 7) ?

OpenStudy (luigi0210):

Yea, not just distribute.

OpenStudy (luigi0210):

*now

OpenStudy (anonymous):

y = 2x - 13?

OpenStudy (luigi0210):

Yup!